OpenStudy (anonymous):

A chemist has two solutions.One contains 30%acid and 70%water and the other contains 70%acid and 30%water.How many litres of each solution should he mix to make 100 litres of solution containing 40%acid and 60% water?

OpenStudy (phi):

say you have x of the 30% and y of the 70% then you can say x+y= 100 also, set up an equation equating the amount of acid: 0.3x + 0.7y = 0.4*100

OpenStudy (phi):

now solve the 2 equations